OpenStudy (unklerhaukus):

Solve the integro-differential equation \[\int\limits_0^xy(u)\,\mathrm du-y'(x)=x\]

OpenStudy (unklerhaukus):

i've got this,

OpenStudy (unklerhaukus):

but the answer in the back of the book says \[=y_0\cosh x-\cosh x+1\\ = (y_0-1)\cosh x+1\]

OpenStudy (zzr0ck3r):

@zzr0ck3r googles integro-differential to see if it is taken for a band name. NOPE!

OpenStudy (kainui):

Maybe this is the problem, but I don't think it should matter so I'll keep looking. \[Y=(y_0-X)\frac{p}{p^2-1} = y_0 \frac{p}{p^2-1} +\frac{1}{1-p^2}\]

OpenStudy (kainui):

Oh nevermind, I think my correction is wrong, I was thinking \(X=\frac{1}{p}\) but really it's \(\frac{1}{p^2}\) isn't it. I guess that's lame, and I suppose you're doing the convolution to avoid partial fractions so I don't blame you.

OpenStudy (kainui):

All of your work looks completely correct to me. Maybe it can be rearranged from one into the other ugh lol.

OpenStudy (unklerhaukus):

sinh = cosh ?

OpenStudy (kainui):

Or your book is wrong. We can try checking to see if the answers satisfy the integro-differential equation.

OpenStudy (kainui):

The book answer is right, your answer is wrong in general but true with \(y_0=0\) so that might give some clues.

ganeshie8 (ganeshie8):

By a different method im getting y(x) = C*coshx + 1

OpenStudy (unklerhaukus):

hmmm

OpenStudy (kainui):

I am not sure but I don't think the inverse laplace transform was done correctly here: \[\mathcal{L}^{-1}\{y_0-X\} = y_0-x\]

ganeshie8 (ganeshie8):

I'll need to review laplace transofrms to understand your method, but your y(x) is clearly wrong as it doesn't satisfy the given equation

OpenStudy (unklerhaukus):

what is the inverse Laplace of a constant?

OpenStudy (kainui):

Dirac delta? Not sure

OpenStudy (unklerhaukus):

oh, yeah the delta thingo

OpenStudy (kainui):

seems like that could plausibly fix it, not gonna work it out though unless you say it doesn't work haha

OpenStudy (unklerhaukus):

how could partial fractions be used instead?

OpenStudy (kainui):

Well if you expand X and avoid the convolution, then you end up with: \[\frac{1}{p^2} \frac{p}{p^2-1} = \frac{1}{p(p^2-1)} = -\frac{1}{p}+\frac{1}{2} \frac{1}{p+1}+\frac{1}{2} \frac{1}{p-1} \] then just transform these terms individually.

OpenStudy (kainui):

Looks like these will inverse transform into some exponentials, which is fine cause we can make cosh and sinh out of exponentials.

OpenStudy (unklerhaukus):

What method did you use @ganeshie8 ?

ganeshie8 (ganeshie8):

Differentiate through out, solve the second order ode, then reduce the constants...

ganeshie8 (ganeshie8):

The solutions of \[\int\limits_0^xy(u)\,\mathrm du-y'(x)=x\tag{1}\] are a subset of the solutions to \[y - y'' = 1\tag{2}\] Not hard to see that \(y = 1\) is a particular solution to \((2)\) The null solution is \(c_1e^x + c_2e^{-x}\). The complete solution to \((2)\) is \(y = 1+c_1e^x + c_2e^{-x}\) Plugging this in \((1)\) gives \(c_1=c_2\), so the general solution to \((1)\) has to be \(y = c(e^x + e^{-x})+1\), which can be rearranged as \(y = C* \cosh x + 1\)

OpenStudy (unklerhaukus):

I can now get your method to work @ganeshie8 , it's not hard to solve 2nd order DE. Thank you for providing it. However I would still like to fix my Laplace convolution method

OpenStudy (unklerhaukus):

I think you are right about my error @Kainui \[\mathcal{L}^{-1}\{y_0-X\} \neq y_0-x\] but i'm not really sure how to fix this

OpenStudy (kainui):

Alright I'll pick it up right where you left it off and give you some stuff to finish it on your own I think cause it should boil down to just distributing mostly the same work you have already done. \[Y= (y_0-X)\frac{p}{p^2-1}\]\[y=\mathcal{L}^{-1}[\mathcal{L}\{y_0\delta(x) -x\}\mathcal{L}\{\cosh(x)\}] = (y_0 \delta(x)-x) \star \cosh(x) \] Actually, if you don't know how to derive the convolution, I think it's definitely worth your time cause it's pretty ubiquitous, it's a way of sidestepping multiplying two series together and looking at only their coefficients and is sorta a meeting point of linear algebra and group theory. Ok so on with the calculation... \[y =\int_0^x [y_0 \delta(x-u) - (x-u)] \cosh(u)du\] One of these integrals you have already solved so I'll just give you what you need: \[\int_0^x \delta(x-u) f(u)du = f(x)\] You can think of \(\delta(0)=\infty\) and all other values \(\delta(t)=0\). It's essentially the identity matrix, and is the identity element of the group of functions with the convolution as the binary operation of this abelian group... Too much information, w/e throw it away lol. That should be enough to solve it now: \[y =y_0 \cosh(x)+\int_0^x - (x-u) \cosh(u)du\]

OpenStudy (kainui):

Yeah continuing this on, I am getting the exact answer as your book now. Awesome.

OpenStudy (kainui):

Just to sorta show the forward way, since x=0 is the only value that \(\delta(x)\) is nonzero, it "pulls out" that value: \[\mathcal{L}\{\delta\}(p)=\int_0^\infty \delta(x) e^{-px} dx = e^{-p0}= 1\]

OpenStudy (unklerhaukus):

\[\boxed{ \newcommand \dd [1] {\,\mathrm d#1} % infinitesimal \newcommand \intl[4] {\int\limits_{#1}^{#2}{#3}{\dd #4}} % integral_a^b{f(x)}dx \newcommand \Lap [1] {\operatorname{\mathcal L}\left\{#1\right\}} % Laplace transform (p) \newcommand \Lin [1] {\operatorname{\mathcal L}^{-1}\left\{#1\right\}} % inverse Laplace transform \newcommand \Ev [3] {\left.{#1}\right|_{#2}^{#3}} % evaluation limits \newcommand \ps [1] {\left(#1\right)} % dynamic parentheses }\] \[ \begin{align} \intl0x{y(u)}u-y'(x) &= x\\ \Lap{1\star y}-\Lap{y'} &= X\\ \frac1p\cdot Y - \ps{pY-y_0} &=\\ \ps{\frac1p-p}Y &= X-y_0\\ Y &= \ps{X-y_0}{\frac1{\frac1p-p}}\\ &= \ps{y_0-X}{\frac p{p^2-1}}\\ y(x)&= \ps{y_0\delta(x)-x}\star\cosh x\\ &= \intl0x{(y_0\delta(x-u)-(x-u))\cosh u}u\\ &= y_0\intl0x{\delta(x-u)\cosh u}u-x\intl0x{\cosh u}u + \intl0x{u\cosh u}u\\ &= y_0\Ev{\cosh u}{}x-x\Ev{\sinh u}0x + \Ev{u\sinh u}0x-\intl0x{\sinh u}u\\ &= y_0\cosh x-x\sinh x + x\sinh x-\Ev{\cosh u}0x\\ &= y_0\cosh x - \cosh x + 1\\ &= \ps{y_0-1}\cosh x+1 \end{align} \]

OpenStudy (unklerhaukus):

We have \[y(x) = (y_0-1)\cosh x+1\] Taking the derivative \[y'(x)=(y_0-1)\sinh x\] and integral: \[\intl0x{y(u)}u= (y_0-1)\intl0x{\cosh u+1}u\\ \qquad\qquad = (y_0-1)\sinh x+x\] Plugging these into the IDE \[\int\limits_0^xy(u)\,\mathrm du-y'(x)=x\\ \]\[LHS = \Big[(y_0-1)\sinh x+x\Big]-\Big[ (y_0-1)\sinh x\Big] \\ \qquad = x\\ \qquad = RHS\]

OpenStudy (irishboy123):

vanilla way, starting at second line \(\dfrac{1}{p} Y - p Y + y_o = \dfrac{1}{p^2}\) \(Y = \dfrac{p}{p^2 -1^2} y_o - \dfrac{1}{p} \dfrac{1}{p^2 - 1^2}\) \( = \mathcal{L} \{ y_o \cosh x\} - \mathcal{L} \{1\} \mathcal{L} \{\sinh x\}\) Second bit is convolution \(y = y_o \cosh x - \int\limits_0^x \sinh \tau d \tau\) \(y = y_o \cosh x - \left. \cosh \tau \right|_0^x\)

OpenStudy (unklerhaukus):

That's much better. @IrishBoy123